OpenStudy (aonz):

HELP PLEASE!!! ABCD is a cyclic quadrilateral in which AB=5. BC=6. CD=7. AD=8. show that cos ADC = 13/43

OpenStudy (aonz):



OpenStudy (aonz):

@kropot72 @terenzreignz @Mertsj Help pls Btw i know angles D+B = 180

OpenStudy (aonz):

@hartnn help please!!!

OpenStudy (mertsj):

There is a formula for this. First let me relabel your picture so you will understand the formula better.

OpenStudy (aonz):

ok... only formula i know is angle B + D=180 and angle A+C = 180

OpenStudy (mertsj):